A monopolist has variable costs of VC = 2q2 and faces a demand curve of P = 30 – q, where P is price and q the quantity sold. (Consider that this demand curve is marginal benefit curve for an individual consumer.) The monopolist engages in first degree price discrimination using a two-part tariff, what is the fixed fee (F) and per-unit fee charged (p)?

(a) F = 16, p = 8

(b) F = 32, p = 16

(c) F = 18, p = 24

(d) F = 8, p = 4

(e) None of the above
Answer #1

From the variable cost function we find MC = 4q

Demand is P = 30 - q.

Under two part tariff, per unit charge is P = MC so we have

30 - q = 4q or q = 30/5 = 6

Hence per unit charge is 4*6 = $24

Fixed fee = CS at this price and quantity = 0.5*(max price - current price)*qty

= 0.5*(30 - 24)*6 = $18

Hence, per unit charge P is $24 and fixed fee is $18.

Select (c) F = 18, p = 24